Fred and Wilma start on a hike at 4 mi/hr. Three hours later barney discovers that they had forgotten their lunch and started after them on his bike at 12 mi/hr. How far had Fred and Wilma traveled when Barney caught up with them?
Answer by mananth(16949)

Fred and Wilma start on a hike at 4 mi/hr. Three hours later barney discovers that they had forgotten their lunch and started after them on his bike at 12 mi/hr. How far had Fred and Wilma traveled when Barney caught up with them?
fred and wilma have traveled 12 miles in three hours when barney starts
Let the distance at which they meet be x
x-12 is the distance fred & wilma travel after barney starts & meets them.
Distance traveled by barney is x
time taken by fred & wilma is x-12/ 3
time taken by barney is x/12
But these two times are same
x-12/3 = x/12
12(x-1
12x-144=3x2)=3x
12x-3x=144
9x=144
x=16
fred and wilma have traveled 12+16 miles when Barney catches up.
x=16